OpenStudy (anonymous):

Someone help! Ok, so I have this equation her that I solved. 2x+3y=1200 3y=-2x+1200 (3/3)y= (-2/3)x+(1200/3) y=(-2/3)x+400 How would I graph this using the slope- intercept method?

OpenStudy (texaschic101):

well..the y intercept is 400. The x intercept, you can find by subbing in 0 for y. 2x + 3(0) = 1200 2x = 1200 x = 600. plot the points (0,400) and (600,0). Then connect the lines. You can even go to the y intercept (0,400) and since the slope is -2/3, you come down 2 and go to the right 3 (thats a point on the line) then go down 2 again and to the right 3...all the way until you cross the x axis at 600.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

@texaschic101 Do you know how I would write the equation in function notation?

OpenStudy (texaschic101):

y = -2/3x + 400 f(x) = -2/3x + 400 ==> function notation